OpenStudy (anonymous):

I need an equation that represents the line that passes through the points (6, –3) and (–4, –9)? Thank in advance!

myininaya (myininaya):

a line has form y=mx+b where m is the slope and b is the y-intercept these are the two things we need to find

myininaya (myininaya):

do you know how to find the slope?

OpenStudy (anonymous):

yes

myininaya (myininaya):

so what is the slope here?

OpenStudy (anonymous):

1,4?

myininaya (myininaya):

\[m=\frac{ changey}{change x}=\frac{-3-(-9)}{6-(-4)}\]

myininaya (myininaya):

thats not what i spelled lol

myininaya (myininaya):

m=(change in y)/(change in x)

OpenStudy (anonymous):

i see

myininaya (myininaya):

\[m=\frac{-3+9}{6+4}=\frac{6}{10}=\frac{3}{5}\]

OpenStudy (anonymous):

ok well that helps a bit

myininaya (myininaya):

so now we can fill in one blank so we have \[y=\frac{3}{5}x+b\] so b is the only thing left to find

myininaya (myininaya):

Use a point given to you in plug in that point to find b

myininaya (myininaya):

like plug in x=6 and y=-3 (this is a point given to you) and now solve for b

myininaya (myininaya):

i must go so maybe someone else can check your answer when you get it

OpenStudy (anonymous):

-5 give or take?

OpenStudy (anonymous):

well thanks...alot

myininaya (myininaya):

\[-3=\frac{3}{5}(6)+b\]

myininaya (myininaya):

\[-3-\frac{18}{5}=b\]

myininaya (myininaya):

\[b=\frac{-3(5)}{5}-\frac{18}{5}=\frac{-15}{5}-\frac{18}{5}=\frac{-23}{5}\]

myininaya (myininaya):

so i go now you have all the pieces of the puzzle now just put them in the right spot in you are done

OpenStudy (anonymous):

ok cool thanks for writing it out rather than just throwin out the answer cuz it helps me learn

OpenStudy (paxpolaris):

Another way of looking at it: If you have 2 points (6, –3) and (–4, –9), you can plug this in the equation y=mx+b to get 2 equations:\[-3=6m+b\]\[-9=-4m+b\] solve the simultaneous equations for m & b